A spinning wheel is slowed down by a brake, giving it a constant angular acceleration of

−4.60 rad/s2.

During a 2.20-s time interval, the wheel rotates through 52.9 rad. What is the angular speed of the wheel at the end of the 2.20-s interval?
rad/s
Answer #1

We can use the equation of motion:

\theta=\theta_{0}+w_{0}\cdot t+a\cdot t^{2}

when   t=0 [s]

\theta=\theta_{0}+w_{0}\cdot (0[s])+(-4.60[\frac{rad}{s^{2}}])\cdot (0[s])^{2}=\theta_{0}

During the 2.20 seg time interval the wheel rotates 52.9 rad, we substitute the values on the equation:

52.9[rad]+\theta_{0}=\theta_{0}+w_{0}\cdot (2.20[s])+(-4.60[\frac{rad}{s^{2}}])\cdot (2.20[s])^{2}

52.9[rad]=w_{0}\cdot (2.20[s])+(-4.60[\frac{rad}{s^{2}}])\cdot (2.20[s])^{2}

Solving for w_{0}

w_{0}=\frac{52.9[rad]+(4.60[\frac{rad}{s^{2}}])\cdot (2.20[s])^{2}}{2.20[s]}=34.16 [\frac{rad}{s}]

Now we use the equation of angular speed:

w=w_{0}+a\cdot t

w=34.16[\frac{rad}{s}]-4.60[\frac{rad}{s^{2}}]\cdot (2.2 [s])=24.04[\frac{rad}{s}]

w=24.04[\frac{rad}{s}]